OpenStudy (anonymous):

You are given a box of 100 silver dollars, all facing heads up. You are instructed to shake the box 2 times; after each shake, you will remove all the dollars that are heads up before shaking again. You may keep all the dollars that are still tails up following the sixth shake. How many dollars will you most likely get to keep?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Can you explain how you got that answer?

OpenStudy (anonymous):

By the way the options are A.6 B.12 C.25 D.50

OpenStudy (anonymous):

In the beginning of the question you said 2 shakes, by the end of the question you said 6 shakes. Which is it?

OpenStudy (anonymous):

probability of a dollar remaining a heads = 1/2 Expected no. of tails remaining therefore after 1st shake = 1/2 * 100 = 50 similaryl after second shake 25

OpenStudy (anonymous):

I am not sure as it is the question that asks for it, not I.

OpenStudy (anonymous):

Well, after each shake, you can expect half of the dollars to be eliminated. So after n shakes the amount of dollars left is expected to be: (1/2)^n times the original amount of dollars. So after two shakes you'll have 1/4 of the dollars left. After six shakes you'll have 1/64 left.

OpenStudy (anonymous):

So do I multiply that with 100 in order to get the answer?

OpenStudy (anonymous):

Kawsam, I do not mean to be rude but you just copy and pasted another persons answer to this same question.

OpenStudy (anonymous):

well how would you know that? because u looked online too! lol. but whatever.. ill delete the comment if u wan me to.

OpenStudy (anonymous):

No, you do not have to do that if you do not want to.

OpenStudy (anonymous):

@Yairenis The question is bad in that it suggests two different amounts of shakes. My guess is that it wants two shakes, which would leave 1/4 of 100 or 25.

OpenStudy (anonymous):

wait.... are u in flvs? @vinnv226

OpenStudy (anonymous):

What is "flvs?"

OpenStudy (anonymous):

nvm! lol its an online virtual school in florida.

OpenStudy (anonymous):

Exactly, what I was going to say.

OpenStudy (anonymous):

Thank you for all your help!